Calculate the mass of argon gas in 40.0 L at STP

1
Expert's answer
2021-11-21T22:40:02-0500

of argon (atomic mass of Ar = 40 ) occupies volume of 22400 cm ^3 at stp


Density of argon = 1.69kg/m^3=1.69kg/m3

Mass =1.69×40=67.6Kg=1.69×40=67.6Kg